OpenStudy (anonymous):

Part 1: Solve the inequality 3(x + 6) >_ 4x - 2 and show all of your work. You can use >= to represent the greater than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. HELP PLEASE

OpenStudy (anonymous):

\[3x+18 \ge 4x-2\]\[-x\ge -20\]\[x\le 20\]

OpenStudy (anonymous):

so that is for part one

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

also were did you get the 20 from please explain

OpenStudy (anonymous):

 To get the 20: 3x-4x >= -2 -18 -x >= -20 x <= 20 (when dividing or multiplying by -1 switch the inequality)